23. A competitive firm can sell its product for a price of $3 in the market (there is a reason the word competitive is underlined and in bold!). Total costs are given below. Fill in the following columns in the table: price, total revenue, marginal revenue, marginal costs, variable costs, fixed costs, profit, and average total cost. (Hint if you get stuck: what are variable costs at a quantity of 0? Therefore, what are fixed costs?).

Quantity Price TR     MR      TC       MC      VC      FC       Profit   ATC

0                                  -         2             -                                            -

1                                              3

2                                              5

3                                              8

4                                              12

5                                              17

How much should the firm produce to maximize profit? What price will it charge? What profit will it make? Is this firm in long-run equilibrium? Why or why not?

Now, graph the MR and MC curves, labeling everything relevant. At what quantity do these curves cross, and what does this tell you? (Be sure to answer this section even if you don’t complete the table, it can be done, and may help you fill it in).
Answer #1

Ans:

Quantity Price TR MR TC MC VC FC Profit ATC
0 3 0 -- 2 -- 0 2 -2 --
1 3 3 3 3 1 1 2 0 3
2 3 6 3 5 2 3 2 1 2.50
3 3 9 3 8 3 6 2 1 2.67
4 3 12 3 12 4 10 2 0 3
5 3 15 3 17 5 15 2 -2 3.40

Explanation:

TR = Price * Quantity

MR = Change in TR / Change in Q

TC = FC + VC

MC = Change in TC / Change in Q

ATC = TC / Q

Fixed costs are available even at zero level of output and remain constant throughout the subsequent level of production.

Ans:  The firm should produce 3 units to maximize profit.

Ans: It will charge $3 .

Explanation:

Under perfect competition , the profit maximizing condition is where price equals marginal cost ( P = MC).

Ans: It will make profit of $1.

Explanation:

Profit = TR - TC = $9 - $8 = $1

Ans: No , this firm is not in long-run equilibrium because at longrun equilibrium , P = MR = MC = ATC

Ans:

MR & MC MR MC 0 1 2 4 5 6 3 Quantity (units)Ans: At quantity of 3 units , these curves cross, and it shows the profit maximizing level of output.